From a lift moving upward with a uniform acceleration 'a', a man throws a ball vertically upwards with a velocity v relative to the lift. The ball comes back to the man after a time t.

Show that a + g = 2 v/t

Grade:Select Grade

3 Answers

Aakash
8 years ago
Take case with respect to the lift
Upward velocity with respect to lift v
As lift is moving upward with a acceleration
acceleration of ball with respect to the lift = acceleration of the ball – acceleration of lift
       =  – g – (a)  =  – (g+a)
using s=ut+(1/2)at2
as it again comes back to same position with respevt to lift therefore displacement is zero
 
0 = vt -(1/2)(a+g)t2
vt=(1/2)(a+g)t2
a+g = 2v/t
 
PLEASE APPROVE IF USEFUL
Mohammad Kavish
8 years ago
Thank u Aakash
Porn
2 years ago
[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ[F62817]∅ _ᵗʰⁱˢ ᵐᵉˢˢᵃᵍᵉ ᶜᵃⁿ'ᵗ ᵇᵉ ᵛⁱᵉʷᵉᵈ ᵇᵉᶜᵃᵘˢᵉ ʸᵒᵘ ᵃʳᵉ ⁿᵒᵒᵇ